# Frequency effects in Linear Discriminative Learning

### v1

@@ Abstract
Word frequency is a strong predictor in most lexical processing tasks (Brysbaert et al.,, 2011). Thus, any model of word recognition needs to account for how word frequency effects arise. The Discriminative Lexicon Model (DLM; Baayen et al., 2018a, ; Baayen et al.,, 2019) models lexical processing with mappings between words' forms and their meanings. Comprehension and production are modelled via linear mappings between the two domains. So far, the mappings within the model can either be obtained incrementally via error-driven learning, a computationally expensive process able to capture frequency effects, or in an efficient, but frequency-agnostic closed-form solution modelling the theoretical endstate of learning (EL) where all words are learned optimally. In the present study we show how an efficient, yet frequency-informed mapping between form and meaning can be obtained (Frequency-informed learning; FIL). We find that FIL well approximates an incremental solution while being computationally much cheaper. FIL shows a relatively low type- and high token-accuracy, demonstrating that the model is able to process most word tokens encountered by speakers in daily life correctly. We use FIL to model reaction times in the Dutch Lexicon Project (Keuleers et al.,, 2010) by means of a Gaussian Location Scale Model and find that FIL predicts well the S-shaped relationship between frequency and the mean of reaction times but underestimates the variance of reaction times for low frequency words. FIL is also better able to account for priming effects in an auditory lexical decision task in Mandarin Chinese (Lee,, 2007), compared to EL. Finally, we used ordered data from CHILDES (Brown,, 1973; Demuth et al.,, 2006) to compare mappings obtained with FIL and incremental learning. We show that the mappings are highly correlated, but that with FIL some nuances based on word ordering effects are lost. Our results show how frequency effects in a learning model can be simulated efficiently by means of a closed-form solution, and raise questions about how to best account for low-frequency words in cognitive models.
@@
Author's Note: This is a preprint. This paper has not been peer reviewed.
@@ 1 Introduction
Word frequency effects are ubiquitous in psycholinguistic research. In fact, word frequency (i.e. the number of times a word occurs in some corpus) is one of the most important predictors in a range of psycholinguistic experiments (Brysbaert et al.,, 2011). In the lexical decision task, where participants are asked to decide whether a presented letter string is a word or not, frequency explains by far the most variance in reaction times, compared to other measures such as neighbourhood density or word length (e.g. Baayen,, 2010; Brysbaert et al.,, 2011): higher frequency words elicit faster reactions (e.g. Rubenstein et al.,, 1970; Balota et al.,, 2004). In word naming, another popular experiment in psycholinguistics where participants have to read aloud presented words, word frequency is less important, but still has a reliable effect: higher frequency words are named faster (e.g. Balota et al.,, 2004). Even though the effect of frequency has long been known and studied, to this day new studies are published confirming the effect in ever larger datasets across different languages (e.g. Balota et al.,, 2007; Keuleers et al.,, 2012, 2010; Aguasvivas et al.,, 2018; Brysbaert et al.,, 2016; Ferrand et al.,, 2010) and proposing new frequency counts, explaining for example individual differences when it comes to the frequency effect (e.g. Brysbaert et al.,, 2018; Kuperman and Van Dyke,, 2013).
Unsurprisingly, a central challenge for models of word recognition has therefore always been to explain word frequency effects. This challenge has been met in various different ways by influential models of word recognition. One of the earliest ideas proposed that words are stored in list-like structures ordered by frequency, such that the most frequent words are found earlier than lower frequency words (e.g. Rubenstein et al.,, 1971). This idea was developed into the ``Search Model of Lexical Access'' by Forster, (1976, 1979). In this model there are ``peripheral access files'' in which words are stored according to their frequency of occurrence. These files hold ``access keys'' to where information about each word is stored in a main file. Words in the access files are grouped into bins based on form characteristics. Thus, the model aims to explain both word frequency effects (words are ordered according to their frequency in the peripheral access files) as well as neighbourhood effects (words with similar form are stored in the same bin). Later iterations of the model also suggested a hybrid model between serial and parallel search, where each bin is searched serially, but all bins are searched in parallel (Forster,, 1994).
@@
The interactive activation model (McClelland and Rumelhart,, 1981; Rumelhart and McClelland,, 1982) is in many ways a successor of the Logogen model. It proposes three levels of representations: one for letter features, one for letters and one for words. There are excitatory and inhibitory connections from letter features to letters and from letters to words. Additionally, there are excitatory and inhibitory connections from words to letters. Finally, there are both excitatory and inhibitory connections within each representational level. The interactive activation model was originally proposed to explain the word superiority effect, i.e. the finding that letters are identified faster within a word than within a random letter string (e.g. Reicher,, 1969). However, the model of course also proposes an account of frequency effects: the resting activation level of the word units are set depending on word frequency, such that high frequency words have a higher resting activation than low frequency words. In this way, Jacobs and Grainger, (1994) were able to show that the interactive activation model shows the same effect of frequency on reaction times in various lexical decision experiments as human participants.
@@
The interactive activation model was followed by the triangle model (Seidenberg and McClelland,, 1989; Harm and Seidenberg,, 2004), which consists of distributed representations for orthography, phonology and semantics, each connected via a set of hidden units. In contrary to the interactive activation model, the weights between layers in this network are not set by the modeller but learned using backpropagation (Rumelhart et al.,, 1986). Thus, the error between the model's prediction and the actual target is reduced each time an input-target pair is encountered. For example, to model reading, the triangle model predicts a word's phonology from its orthography. Then, the error between the predicted phonology and the correct phonology of the word is computed, and the model's weights are updated such that the next time the same phonology is to be produced from the same orthography, it will be more accurate. This means that words that are presented to the model more frequently will over time produce the lowest prediction error and are thus recognised faster and more accurately. Therefore, word frequency effects arise not as a consequence of manually changing resting activation levels but from the weights within the network changing according to the input distribution.
@@
A final model of word recognition is the Bayesian Reader Model (Norris,, 2006). This model not only accounts for the frequency effect but also aims to explain why frequency effects should arise in the first place. The model is a simple Bayesian model that integrates a word's prior probability (for which Norris,, 2006, uses its frequency) with the incoming evidence. Thus, high frequency words are recognised faster than low frequency ones. According to Norris, (2006, 2013) this constitutes an ``ideal observer'' model, solving the task at hand as optimally as possible. This explains not only why frequency effects should arise in the first place but also why they play out differently in different experiments.
@@
Interestingly, reaction times for example in lexical decision are best predicted not by raw word frequencies but by log- or rank-transformed frequencies. Again, the various models account for this in different ways. Since the serial search model incorporates word lists, it directly predicts a rank transformation of word frequencies (see also Murray and Forster,, 2004). In the interactive activation model, the resting activation levels of the word units are set according to the words' log frequency (McClelland and Rumelhart,, 1989, Chapter 7). In the triangle model, training items are sampled such that they have a probability proportional to $\log(\text{frequency}+2)$. Note that Seidenberg and McClelland, (1989) do this for practical rather than principled reasons, as training their model on the full token frequencies was too computationally expensive. Finally, the Bayesian Reader utilises raw word frequencies.
@@
We now turn to a newer model of word comprehension and production, the Discriminative Lexicon Model (DLM; Baayen et al., 2018a, ; Baayen et al.,, 2019). This model provides a perspective on the mental lexicon in which mappings between numeric representations for form and meaning play a central role. This model conceptualizes comprehension as involving mappings from high-dimensional modality-specific form vectors to high-dimensional representations of meaning. The initial stage of speech production is modeled as involving a mapping in the opposite direction, starting with a high-dimensional semantic vector (known as embeddings in computational linguistics) and targeting a vector specifying which phone combinations drive articulation. The DLM has been successful in modelling a range of different morphological systems (e.g. Chuang et al.,, 2020, 2022; Denistia and Baayen,, 2021; Nieder et al.,, 2021; Heitmeier et al.,, 2021) as well as behavioural data such as acoustic durations (Schmitz et al.,, 2021; Stein and Plag,, 2021; Chuang et al.,, 2022), (primed) lexical decision reaction times (Gahl and Baayen,, 2023; Heitmeier et al.,, 2022) and data from patients with aphasia (Heitmeier and Baayen,, 2020).
@@
Thus, the DLM proposes that frequency effects arise due to the distribution of the input data: higher frequency words occur more often in the input data, and therefore, the prediction error will be smallest for high frequency words (see Chuang et al.,, 2021; Heitmeier et al.,, 2021, for studies utilising WHL to obtain frequency-informed mapping matrices). Word frequencies are not stored explicitly; rather, they have effects on the weights in the mappings. This account is similar to how frequency effects arise in the triangle model. Similar to the triangle model, the DLM also suffers from computational issues: training on the full frequency distribution with WHL is computationally very demanding. However, the DLM has an advantage over the triangle model: EL only involves a single linear mapping for which a closed-form solution exists. In the present study we introduce a new method which takes frequency of use into account but is computationally efficient by making use of a closed-form solution (``Frequency-informed learning'', FIL). FIL can be used instead of the already established WHL and EL to compute mapping matrices in the DLM.
In the following we compare the three different methods of estimating mapping matrices in the DLM. We show how the model is able to account for frequency effects using WHL and FIL. We demonstrate that FIL is equivalent to training the model incrementally on full token frequencies and is superior to utilising log-transformed frequencies. We show how the DLM is able to model reaction times linearly without the need of log- or rank-transformations. Finally, we investigate what role the order in which words are learned plays in word recognition.
This study is structured as follows: we first lay out the basic concepts underlying linear discriminative learning. Next, we discuss EL, and subsequently, WHL. Against this background, we proceed with proposing a new method computing non-incremental, yet frequency-informed learning (FIL). We then present three case studies, one using FIL to model visual lexical decision latencies from the Dutch Lexicon Project (Keuleers et al.,, 2010), one where we use FIL to model spoken word recognition in Mandarin and a third where we compare WHL and FIL in modelling first word learning with data from CHILDES (Brown,, 1973; Demuth et al.,, 2006). A discussion section brings this study to a close.
@@ 2 Linear Discriminative Learning: Basic concepts and notation
$\mathbf{SG}$ $=$ $\mathbf{C}.$
Given $\mathbf{F}$ and $\mathbf{G}$, we can estimate the predicted semantic vectors
@@
To evaluate the accuracy of a comprehension mapping, the predicted semantic row vectors $\hat{\mathbf{s}}$ of $\hat{\mathbf{S}}$ are correlated with all the corresponding semantic row vectors in the gold standard semantic matrix $\mathbf{S}$. If the predicted semantic vector of a word form is closest to its target vector, it is counted as correct. This accuracy measure is referred to as correlation accuracy in the following, and we will sometimes denote it as accuracy@$1$As only 2.7% of the words in our Dutch dataset are homographs, for simplicity we use ‘strict evaluation’ throughout this study, i.e. we do not take homographs into account while evaluating accuracy (further information on the various methods to evaluate accuracy in Heitmeier et al.,, 2021).. More lenient accuracy measures accuracy@k accept model performance as satisfying when $\hat{\mathbf{s}}$ is among the top $k$ nearest semantic vectors of the gold standard semantic vector $\mathbf{s}$. For detailed introductions to the DLM, see Baayen et al., 2018a ; Baayen et al., (2019); Heitmeier et al., (2021, 2023).
3 Data
@@
For the present study, we used a subset of 2646 singular and plural Dutch nouns and verbs (for which frequency was at least 1) taken from a dataset originally extracted from the Dutch CELEX database (Baayen et al.,, 1995) by Ernestus and Baayen, (2003) [^2] Data, code and statistical models presented in this study are available at https://osf.io/h2szj/.. These words have monomorphemic stems ending in an obstruent that is realized as voiceless when word-final but that in syllable onset appears with voicing in some words and without voicing in others. We used 300-dimensional Dutch fasttext embeddings (Grave et al.,, 2018) to represent semantics [^3] https://fasttext.cc/docs/en/crawl-vectors.html#models. Since we were unable to obtain word embeddings for all words in our dataset, this left us with 2638 word forms (i.e. excluding 8 word forms). Because this dataset is quite small, we also present results with a larger dataset extracted from the Dutch Lexicon Project (Keuleers et al.,, 2010) in later sections. The frequencies of use that we use in this study are taken from CELEX. In what follows, we first present results that make use of bi-grams for representing word forms (used in their orthographic representation), resulting in a dimensionality of 360. The use of bigrams is motivated by the choice to minimize the carbon footprint of our simulations. However, for the new, computationally highly efficient frequency-sensitive model that is at the heart of this study, we made use of trigrams.
In what follows, for expositional simplicity, we focus mainly on comprehension. We note, however, that frequency-informed mappings are equally important for modeling production.
4 Endstate Learning

### v2

@@ Abstract
Word frequency is a strong predictor in most lexical processing tasks (Brysbaert et al.,, 2011). Thus, any model of word recognition needs to account for how word frequency effects arise. The Discriminative Lexicon Model (DLM; Baayen et al., 2018a, ; Baayen et al.,, 2019) models lexical processing with mappings between words' forms and their meanings. Comprehension and production are modelled via linear mappings between the two domains. So far, the mappings within the model can either be obtained incrementally via error-driven learning, a computationally expensive process able to capture frequency effects, or in an efficient, but frequency-agnostic solution modelling the theoretical endstate of learning (EL) where all words are learned optimally. In the present study we show how an efficient, yet frequency-informed mapping between form and meaning can be obtained (Frequency-informed learning; FIL). We find that FIL well approximates an incremental solution while being computationally much cheaper. FIL shows a relatively low type- and high token-accuracy, demonstrating that the model is able to process most word tokens encountered by speakers in daily life correctly. We use FIL to model reaction times in the Dutch Lexicon Project (Keuleers et al.,, 2010) by means of a Gaussian Location Scale Model and find that FIL predicts well the S-shaped relationship between frequency and the mean of reaction times but underestimates the variance of reaction times for low frequency words. FIL is also better able to account for priming effects in an auditory lexical decision task in Mandarin Chinese (Lee,, 2007), compared to EL. Finally, we used ordered data from CHILDES (Brown,, 1973; Demuth et al.,, 2006) to compare mappings obtained with FIL and incremental learning. We show that the mappings are highly correlated, but that with FIL some nuances based on word ordering effects are lost. Our results show how frequency effects in a learning model can be simulated efficiently, and raise questions about how to best account for low-frequency words in cognitive models.
@@ 1 Introduction
Word frequency effects are ubiquitous in psycholinguistic research. In fact, word frequency (i.e. the number of times a word occurs in some corpus) is one of the most important predictors in a range of psycholinguistic experimental paradigms (Brysbaert et al.,, 2011). In the lexical decision task, where participants are asked to decide whether a presented letter string is a word or not, frequency explains by far the most variance in reaction times, compared to other measures such as neighbourhood density or word length (e.g. Baayen,, 2010; Brysbaert et al.,, 2011): higher frequency words elicit faster reactions (e.g. Rubenstein et al.,, 1970; Balota et al.,, 2004). In word naming, another popular experimental paradigm in psycholinguistics where participants have to read aloud presented words, word frequency is less important, but still has a reliable effect: higher frequency words are named faster (e.g. Balota et al.,, 2004). Even though the effect of frequency has long been known and studied, to this day new studies are published confirming the effect in ever larger datasets across different languages (e.g. Balota et al.,, 2004; Keuleers et al.,, 2012, 2010; Brysbaert et al.,, 2016; Ferrand et al.,, 2010). Some studies have also proposed new frequency counts, explaining for example individual differences when it comes to the frequency effect (e.g. Brysbaert et al.,, 2018; Kuperman and Van Dyke,, 2013).
A central challenge for models of word recognition is therefore to explain word frequency effects. This challenge has been met in various different ways by influential models of word recognition. One of the earliest ideas proposed that words are stored in list-like structures ordered by frequency, such that the most frequent words are found earlier than lower frequency words (e.g. Rubenstein et al.,, 1971). This idea was developed into the ``Search Model of Lexical Access'' by Forster, (1976, 1979). In this model there are ``peripheral access files'' in which words are stored according to their frequency of occurrence. These files hold ``access keys'' to where information about each word is stored in a main file. Words in the access files are grouped into bins based on form characteristics. Thus, the model aims to explain both word frequency effects (words are ordered according to their frequency in the peripheral access files) as well as neighbourhood effects (words with similar form are stored in the same bin). Later iterations of the model also suggested a hybrid model between serial and parallel search, where each bin is searched serially, but all bins are searched in parallel (Forster,, 1994).
@@
The interactive activation model (McClelland and Rumelhart,, 1981; Rumelhart and McClelland,, 1982) is in many ways a successor of the Logogen model. It proposes three levels of representations: one for letter features, one for letters and one for words. There are excitatory and inhibitory connections from letter features to letters and from letters to words. Additionally, there are excitatory and inhibitory connections from words to letters. Finally, there are both excitatory and inhibitory connections within each representational level (note though that the feature-feature inhibition was set to zero in the original implementation). The interactive activation model was originally proposed to explain the word superiority effect, i.e. the finding that letters are identified faster within a word than within a random letter string (e.g. Reicher,, 1969). However, the model also proposes an account of frequency effects: the resting activation level of the word units are set depending on word frequency, such that high frequency words have a higher resting activation than low frequency words. In this way, Jacobs and Grainger, (1994) were able to show that the interactive activation model shows the same effect of frequency on reaction times in various lexical decision experiments as human participants.
@@
The interactive activation model was followed by the triangle model (Seidenberg and McClelland,, 1989; Harm and Seidenberg,, 2004), which consists of distributed representations for orthography, phonology and semantics, each connected via a set of hidden units. In contrast to the interactive activation model, the weights between layers in this network are not set by the modeller but learned using backpropagation (Rumelhart et al.,, 1986). Thus, the error between the model's prediction and the actual target is reduced each time an input-target pair is encountered. For example, to model reading, the triangle model takes as input a word's orthography and predicts its phonology. Then, the error between the predicted phonology and the correct phonology of the word is computed, and the model's weights are updated such that the next time the same phonology is to be produced from the same orthography, it will be more accurate. The more often a word is presented to the model, the more accurate its predicted phonology becomes. This means that high frequency words will over time produce the lowest prediction error and are thus recognised faster and more accurately. Therefore, word frequency effects arise not as a consequence of manually changing resting activation levels but from the weights within the network changing according to the input distribution.
@@
A final model of word recognition reviewed here is the Bayesian Reader Model (Norris,, 2006). This model not only accounts for the frequency effect but also aims to explain why frequency effects should arise in the first place. The model is a simple Bayesian model that integrates a word's prior probability (for which Norris,, 2006, uses its frequency) with the incoming evidence. Thus, high frequency words are recognised faster than low frequency ones. According to Norris, (2006, 2013) this constitutes an ``ideal observer'' model, solving the task at hand as optimally as possible. This explains not only why frequency effects should arise in the first place but also why they play out differently in different experiments.
@@
Interestingly, reaction times for example in lexical decision are best predicted not by raw word frequencies but by log- or rank-transformed frequencies. Again, the various models account for this in different ways. Since the serial search model incorporates word lists, it directly predicts a rank transformation of word frequencies (see also Murray and Forster,, 2004). In the interactive activation model, the resting activation levels of the word units are set according to the words' log frequency (McClelland and Rumelhart,, 1989, Chapter 7). In the triangle model, training items are sampled such that they have a probability proportional to $\log(\text{frequency}+2)$. Note that Seidenberg and McClelland, (1989) do this for practical rather than principled reasons, as sampling training items proportionally to full token frequencies would require orders of magnitude more computation to achieve the same coverage (this is a practice adopted also in other work, see for instance Li et al.,, 2007). Finally, the Bayesian Reader utilises raw word frequencies.
@@
We now turn to a more recent model of word comprehension and production, the Discriminative Lexicon Model (DLM; Baayen et al., 2018b, ; Baayen et al.,, 2019). This model provides a perspective on the mental lexicon in which mappings between numeric representations for form and meaning play a central role. This model conceptualizes comprehension as involving mappings from high-dimensional modality-specific form vectors to high-dimensional representations of meaning. The initial stage of speech production is modeled as involving a mapping in the opposite direction, starting with a high-dimensional semantic vector (known as embeddings in computational linguistics) and targeting a vector specifying which phone combinations drive articulation. The DLM has been successful in modelling a range of different morphological systems (e.g. Chuang et al.,, 2020, 2022; Denistia and Baayen,, 2021; Nieder et al.,, 2023; Heitmeier et al.,, 2021) as well as behavioural data such as acoustic durations (Schmitz et al.,, 2021; Stein and Plag,, 2021; Chuang et al.,, 2022), (primed) lexical decision reaction times (Gahl and Baayen,, 2023; Heitmeier et al., 2023a, ) and data from patients with aphasia (Heitmeier and Baayen,, 2020).
@@
Thus, the DLM proposes that frequency effects arise due to the distribution of the input data: higher frequency words occur more often in the input data, and therefore, the prediction error will be smallest for high frequency words (see Chuang et al.,, 2021; Heitmeier et al.,, 2021, for studies utilising WHL to obtain frequency-informed mapping matrices). Word frequencies are not stored explicitly; rather, they have effects on the weights in the mappings. This account is similar to how frequency effects arise in the triangle model. Similar to the triangle model, the DLM also suffers from computational issues: training on the full frequency distribution with WHL is computationally very demanding.
Recent modelling efforts with the DLM have been limited by the disadvantages of EL and WHL: they either had to opt for EL, which resulted in models that were not informed about word frequencies (e.g. Heitmeier et al., 2023a, ), or for WHL, which limited the amount of data the models could be trained on (e.g. Heitmeier et al.,, 2021; Chuang et al.,, 2021). The present paper aims to solve this problem by introducing a new method for computing the mapping matrices that takes frequency of use into account but is computationally efficient by making use of a numerically efficiently solvable solution: ``Frequency-informed learning'' (FIL). FIL can be used instead of the already established WHL and EL to compute mapping matrices in the DLM.
In the following we compare the three different methods of estimating mapping matrices in the DLM. We show how the model is able to account for frequency effects using WHL and the newly introduced FIL. We demonstrate that FIL is equivalent to training the model incrementally on full token frequencies and is superior to utilising log-transformed frequencies. We show how the DLM is able to model reaction times linearly without the need of log- or rank-transformations. Finally, we investigate what role the order in which words are learned plays in word recognition.
This study is structured as follows: we first lay out the basic concepts underlying linear discriminative learning in Section 2. In Section 3, we discuss EL, and subsequently, WHL. Against this background, we proceed with proposing a new method computing non-incremental, yet frequency-informed learning (FIL). We then present three case studies, one using FIL to model visual lexical decision latencies from the Dutch Lexicon Project (Keuleers et al.,, 2010) (Section 4.1), one where we use FIL to model spoken word recognition in Mandarin (Section 4.2) and a third where we compare WHL and FIL in modelling first word learning with data from CHILDES (Brown,, 1973; Demuth et al.,, 2006) (Section 5). A discussion section brings this study to a close.
@@ 2 Linear Discriminative Learning: Basic concepts and notation
$\mathbf{SG}$ $=$ $\mathbf{C}$
where $\mathbf{CF}$ and $\mathbf{SG}$ refer to the matrix multiplication of $\mathbf{C}$ and $\mathbf{F}$ and of $\mathbf{S}$ and $\mathbf{G}$ respectively. Further information on this operation can be found for instance in Beaumont, (1965). Given $\mathbf{F}$ and $\mathbf{G}$, we can estimate the predicted semantic vectors
@@
To evaluate the accuracy of a comprehension mapping, the predicted semantic row vectors $\hat{\mathbf{s}}$ of $\hat{\mathbf{S}}$ are correlated with all the corresponding semantic row vectors in the gold standard semantic matrix $\mathbf{S}$. If the predicted semantic vector of a word form is closest to its target vector, it is counted as correct. This accuracy measure is referred to as correlation accuracy in the following, and we will sometimes denote it as accuracy@$1$As only 2.7% of the words in our Dutch dataset are homographs, for simplicity we use ‘strict evaluation’ throughout this study, i.e. we do not take homographs into account while evaluating accuracy (further information on the various methods to evaluate accuracy in Heitmeier et al.,, 2021).. More lenient accuracy measures accuracy@k accept model performance as satisfying when s is among the top k nearest target semantic vectors of the predicted semantic vector $\hat{\mathbf{s}}$. For detailed introductions to the DLM, see Baayen et al., 2018a ; Baayen et al., (2019); Heitmeier et al., (2021); Heitmeier et al., 2023b .
3 Three methods for computing mappings in the DLM
This section introduces the two existing methods for computing mappings in the DLM, Endstate Learning (EL) and Widrow-Hoff learning (WHL), and explains their respective disadvantages using a small Dutch dataset from Ernestus and Baayen, (2003). Finally, we present Frequency-informed learning (FIL). Since it is much more computationally efficient than WHL, we also demonstrate its usage on a larger dataset from Keuleers et al., (2010). For expositional simplicity, we focus mainly on comprehension. We note, however, that frequency-informed mappings are equally important for modeling production.
3.1 Setup
3.1.1 Data
For the present section, we used two datasets:
@@ 3.1.1 Data
Small dataset: A subset of 2,646 singular and plural Dutch nouns and verbs (for which frequency was at least 1) taken from a dataset originally extracted from the Dutch CELEX database (Baayen et al.,, 1995) by Ernestus and Baayen, (2003) [^2] Data, code and statistical models presented in this study are available at https://osf.io/h2szj/. These words have monomorphemic stems ending in an obstruent that is realized as voiceless when word-final but that in syllable onset appears with voicing in some words and without voicing in others. We used 300-dimensional Dutch fasttext embeddings (Grave et al.,, 2018) to represent semantics [^3] https://fasttext.cc/docs/en/crawl-vectors.html#models. Since we were unable to obtain word embeddings for all words in our dataset, this left us with 2,638 word forms (i.e. excluding 8 word forms).
Large dataset: We also present results with a larger dataset extracted from the Dutch Lexicon Project (DLP, Keuleers et al.,, 2010) in later sections. We used all 13,669 words from the DLP for which we were able to obtain fasttext embeddings.
The frequencies of use that we use in this study when working both with the small and the large datasets are taken from CELEX.
3.1.2 Modelling choices
In what follows, we present results with two different form vector setups:
Low-dimensional form vectors: We make use of bi-grams for representing word forms (used in their orthographic representation), resulting in a dimensionality of 360. The use of bigrams is motivated by the choice to minimize the carbon footprint and duration of our simulations, especially when using WHL.
High-dimensional form vectors: For the new, computationally highly efficient frequency-sensitive model that is at the heart of this study, we made use of trigrams. For the small dataset this resulted in a form vector dimensionality of 1,776, and 4,678 for the large dataset.
3.2 Endstate Learning
The first implementation of the DLM in the R (R Core Team,, 2020) package WpmWithLdl (Baayen et al., 2018b, ) estimates the `endstate' of learning. This implementation constructs a mapping $\mathbf{F}$ between the cue matrix $\mathbf{C}$ and the semantic matrix $\mathbf{S}$ using the pseudo-inverse by solving the following set of equations:
$\mathbf{CF}$ $=\mathbf{S}$
$\mathbf{C}^{T}\mathbf{CF}$ $=\mathbf{C}^{T}\mathbf{S}$
$(\mathbf{C}^{T}\mathbf{C})^{-1}\mathbf{C}^{T}\mathbf{CF}$ $=(\mathbf{C}^{T}\mathbf{C})^{-1}\mathbf{C}^{T}\mathbf{S}$
$\mathbf{IF}$ $=(\mathbf{C}^{T}\mathbf{C})^{-1}\mathbf{C}^{T}\mathbf{S}$
$\mathbf{F}$ $=(\mathbf{C}^{T}\mathbf{C})^{-1}\mathbf{C}^{T}\mathbf{S}$ (1)
where $\mathbf{I}$ denotes the identity matrix and $\mathbf{C}^{T}$ the transpose of $\mathbf{C}$. Details on these equations (known as the normal equations in statistics) can be found in Faraway, (2005).